SpaceRef

SpaceRef


NASA Hubble Space Telescope Daily Report #4206

Status Report From: Space Telescope Science Institute
Posted: Wednesday, September 27, 2006

image

HUBBLE SPACE TELESCOPE DAILY REPORT # 4206

- Continuing to collect World Class Science

PERIOD COVERED: UT September 25, 2006 (DOY 268)

OBSERVATIONS SCHEDULED

NIC1/NIC2/NIC3 8794

NICMOS Post-SAA calibration - CR Persistence Part 5

A new procedure proposed to alleviate the CR-persistence problem of NICMOS. Dark frames will be obtained immediately upon exiting the SAA contour 23, and every time a NICMOS exposure is scheduled within 50 minutes of coming out of the SAA. The darks will be obtained in parallel in all three NICMOS Cameras. The POST-SAA darks will be non-standard reference files available to users with a USEAFTER date/time mark. The keyword 'USEAFTER=date/time' will also be added to the header of each POST-SAA DARK frame. The keyword must be populated with the time, in addition to the date, because HST crosses the SAA ~8 times per day so each POST-SAA DARK will need to have the appropriate time specified, for users to identify the ones they need. Both the raw and processed images will be archived as POST-SAA DARKSs. Generally we expect that all NICMOS science/calibration observations started within 50 minutes of leaving an SAA will need such maps to remove the CR persistence from the science images. Each observation will need its own CRMAP, as different SAA passages leave different imprints on the NICMOS detectors.

NIC2/NIC1 11000

Evaporating Disks

Evaporation in the vicinity of an O star disrupts protoplanetary disks, as seen in the Orion proplyds. We have found a number of evaporating disks with Spitzer, which are in some ways more dramatic and better oriented for detailed study than the proplyds ? they have cometary tails extending up to 0.1 pc from the evaporation working surface. We will use Spitzer/IRS and HST/NICMOS to investigate these systems in more detail. We want to explore the excitation condition in the gas, both in the head and in the tail where possible. We will measure the effects of evaporation on the characteristic emission features of the dust. We also will use NICMOS to image them in detail, including mapping complex structures resolved in their tails at 24 microns

NIC3 10780

The Unusual Afterglow and Host Galaxy of the Short GRB 060121

We request Director's Discretionary Time to observe the afterglow and host galaxy of the short- hard gamma-ray bursts GRB 060121. This is only the fourth time an optical afterglow of a short- hard burst has been found at its properties are significantly different to other cases. Both the afterglow and host galaxy are much fainter than previous short bursts {for which optical afterglows have been located} and the afterglow may also show the signature of dust extinction. Such extinction is completely unexpected for short bursts, given the currently popular model of their origin in NS-NS mergers. We propose ACS and NICMOS observations which will locate the afterglow and host galaxy of GRB 060121, provide constraints on its beaming angle, search for the signature of dust, and measure the offset from its host. These observations offer a strong test of NS-NS merger models.

WFPC2 10748

WFPC2 CYCLE 14 Standard Darks

This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. Over an extended period these data will also provide a monitor of radiation damage to the CCDs.

WFPC2 10822

CIII] Imagery of Planetary Nebulae

We propose to image five planetary nebulae {PNe} with the F185W filter of WFPC2 in order to study the spatial distribution of the ultraviolet C III] 1909 Angstrom line relative to [O III] 5007 Angstrom and other optical lines. This program follows a Cycle 12 SNAP program {GO 9740} of WFPC2 F185W imaging of PNe and H II regions which validated the feasibility of obtaining images of bright PNe in C III] through careful continuum subtraction and calibrations based on archival IUE SWP and LWP/R spectra. However, in the SNAP program only short {10 minute} exposures of four PNe were taken; we now propose much longer exposures {60-72 minutes} to obtain the desired signal-to-noise to develop high quality C++ ionization maps for comparison with O++ and other ions via photoionization modeling. The five PNe chosen: NGC 2392, NGC 3242, NGC 6543, NGC 6720, and NGC 7662 were selected on the basis of their high surface brightness, extensive UV spectra available from IUE and HST, and numerous WFPC2 images in the archives in other emission lines. We expect the results of this study to impact the utility and accuracy of using the C III] 1909 lines compared to [O III] 5007 for C/O abundance derivations in PNe and H II regions.

FLIGHT OPERATIONS SUMMARY:

Significant Spacecraft Anomalies: (The following are preliminary reports of potential non-nominal performance that will be investigated.)

HSTARS: (None)

COMPLETED OPS REQUEST: (None)

COMPLETED OPS NOTES: (None)
                          SCHEDULED      SUCCESSFUL 
FGS GSacq                04                     04 
FGS REacq                04                     04 
OBAD with Maneuver   16                    16 

SIGNIFICANT EVENTS:

Flash Report: Update on ACS DOY 266 Suspend (by R. Stevens, Mon, 25 Sep 14:06:21) The Tiger Team provided an update to the ACS Suspend event that occurred on day 266. The briefing was held in GSFC Building 3, S107A at 9:00 am Eastern. The tiger team reported that additional analysis confirmed that the +35 volt power was not getting to the ASPC2 electronics of the HRC CEB. This was not a telemetry monitoring problem. To better understand the possible recovery options, the Tiger Team has recommended NSSC-1 reconfiguration to get better insight in to the relay positions . Currently, the K2 relay status is sampled as a bilevel telemetry. The NSSC-I will be patched to collect the status as an analog voltage. The voltage will be compared to the PSPICE model to determine if the K2 relay is pointing to LVPS1, LVPS2, or some indeterminate value. Data collected after the patch along with hardware relay info from the GSFC Parts Branch and stress analysis of the HRC detector will help determine the type of recovery commanding required to get ACS operational in the near-term. HST Project gave the go ahead to proceed with preparations for the patch. Testing will proceed in the ESTIF today. The next status meeting will be an Ops Briefing on the ACS Relay Monitoring Test. This is scheduled for Tuesday, September 26 at 9:00 am Eastern in S107A. A new website has been established to support Tiger Team activities http://edocs.hst.nasa.gov/hstsysman/ACSHRCARB/default.aspx

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter